OpenStudy (anonymous):

A survey conducted in a certain town found that 1 out of every 5 individuals owned a bike. A statistician calculated the margin of error to be 1.5%. What is the maximum percentage of individuals who own a bike? 21.5% 20% 6.5% 30%

OpenStudy (anonymous):

I think it is 20% because I did \[\frac{ 1 }{ 5 }\times 100= 0.02 \times 100 = 20%\]

Vocaloid (vocaloid):

1/5 individuals means that 20% of the sample owns a bike margin of error of 1.5% means that the actual value could be between 20-1.5% to 20+1.5% which means that the minimum value is 18.5% and the maximum value is 21.5%

OpenStudy (anonymous):

A is correct, thank youu